OpenStudy (anonymous):

A salesperson receives a weekly salary of $250 plus a commission of $12 for each computer sold,n. Weekly pay is given by the equation P=12n+250.Graph the equation for n between 0 and 50 computers, and explain the meaning of the slope and y-intercept. P.S. I already explained the meaning of the slope and y-intercept,But I don't know how I'm supposed to graph this equation.

OpenStudy (anonymous):

In this equation n will be the x coordinate and p will be the y coordinate. So find the p coordinate when n=0 and n=50 then draw a line through them

OpenStudy (anonymous):

How in the world do I find the p coordinate? ._.

OpenStudy (anonymous):

Plug in the n value and then solve the equation for p.

OpenStudy (anonymous):

So you wanna know the coordinate for 0 computers. (0, ?) p=12(0)+250 p=250 so the point is (0, 250)

OpenStudy (anonymous):

Do the same for 50

OpenStudy (anonymous):

So it's p=12(50+250 p=750+250 and then I divide?

OpenStudy (anonymous):

p=12(50)+250 p=600+250 p=850

OpenStudy (anonymous):

My bad again.Stupid "scientific calculator"

OpenStudy (anonymous):

Ha yeah. Now plot (0, 250) and (50, 850) and draw a straight line through them